OpenStudy (anonymous):

.Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? A) 3 adults and 12 kids B) 5 adults and 6 kids C) 7 adults and 4 kids D) 8 adults and 2 kids

jimthompson5910 (jim_thompson5910):

Choice A says " 3 adults and 12 kids " So that means a = 3, k = 12

jimthompson5910 (jim_thompson5910):

Plug a = 3, k = 12 into the equation to get 8a + 4k = 72 8*3 + 4k = 72 ... replace 'a' with 3, since a = 3 8*3 + 4*12 = 72 ... replace k with 12, because k = 12 24 + 48 = 72 72 = 72 ... this equation is TRUE, so a = 3 and k = 12 both together form an ordered pair solution (3,12) to the equation 8a + 4k = 72

jimthompson5910 (jim_thompson5910):

So "A) 3 adults and 12 kids " is a possible combination That means we eliminate choice A because they ask " Which is NOT a possible combination of adults and kids?"

jimthompson5910 (jim_thompson5910):

Hopefully you're seeing how this all works?